# Do spectral cues matter in contrast-based graph self-supervised learning?

### v1

@@ Appendix E More experiments
E.1 Graph-Level analysis for degeneration of the spectrum after EP (Sec. 7.1 Cont.)
For graph-level analysis, we basically follow the settings mentioned above in node-level one. The only difference from the node-level task is that we have multiple original graphs with various numbers of nodes, leading to the inconsistent dimensions of the vector of the eigenvalues. Therefore, to provide a more detailed comparison of spectral properties at the graph level, we employ Kernel Density Estimation (KDE) [23] to interpolate and smooth the distributions of eigenvalues. We compare two groups of graph spectra. Each group’s spectra are processed to compute their KDEs, and the mean and standard deviation of these KDEs are calculated.
Like the results in node-level analysis, in Fig. 4(a) and 4(b), we witness the obvious difference between the average spectra of original graphs while the significant overlap between those of augmented graphs, especially if pay attention to the overlapping of the area created by the standard deviation of KDEs. Again, this contrast is not trivial because of the striking mismatch between the average spectra of original and augmented graphs in both datasets, as presented in Fig. 4(c) and 4(d).
(a) Average spectra of original graphs
(b) Average spectra of augmented graphs
(c) The Spectrum of MUTAG
(d) The Spectrum of PROTEINS
E.2 Relationship between spectral cues and performance of EP
@@
Based on the findings obtained from Sec 7.1, it is very likely that spectral information can not be distinguishable enough for good representation learning on the graph. But to more directly answer the question of whether spectral cues play an important role in the learning performance of EP, we continue to conduct a statistical analysis to evaluate the influence of various factors on the learning performance. The results turn out to be consistent with our claim that spectral cues are insignificant aspects of outstanding performance on accuracy observed in Sec. 6.
E.2.1 Statistical analyses on key factors on performance of EP
@@
From a statistical angle, we have a few dimensions of factors that can possibly influence learning performance, like the parameters of EP (i.e. drop rate $p$ in DropEdge or add rate $q$ in AddEdge) as well as potential spectral cues lying in the argument graphs. Therefore, to rule out the possibility that spectral cues are significant, comparisons are conducted on the impact of the parameters of EP in the augmentations versus:
@@
The polynomial regression analysis in Table 6 highlights that the drop rate $q$ is the primary factor influencing model performance, showing strong and significant linear and non-linear relationships with test accuracy. In contrast, both the OG-AUG and AUG-AUG spectral distances have relatively minor impacts on performance, indicating that they are not significant determinants of the model’s efficacy.
@@
Table 7: IV2SLS regression results for the node-level task. The parameter $q$ refers to the drop rate in DropEdge. The experiment comes in pairs for each pair of variables and the better result is marked in grey.
@@
(X = AUG-AUG, Z = $q$) 0.341 45.77 1.68e-08
(Z = $q$ ,Z = AUG-AUG) 0.611 47.85 9.85e-09
(X = OG-AUG, Z = $q$) 0.250 40.22 7.51e-08
(X = $q$, Z = OG-AUG) 0.606 41.27 5.62e-08
The IV2SLS regression results for the node-level task in Table 7 indicate that the edge perturbation parameters are more significant determinants of model performance than spectral distances. Specifically, when the spectral distance between augmented graphs (AUG-AUG) is the explanatory variable (X) and drop rate $q$ are the instrument (Z), the model explains 34.1% of the variance in performance (R-squared = 0.341). Conversely, when the roles are reversed (X = $q$, Z = AUG-AUG), the model explains 61.1% of the variance (R-squared = 0.611), indicating a stronger influence of edge perturbation parameter $q$. A similar conclusion can be made when comparing OG-AUG and $q$.
@@
The analyses distinctly show that the direct edge perturbation parameters have a consistently stronger and more significant impact on model performance than the two types of spectral distances that serve as a reflection of spectral cues. The results support the argument that while spectral cues might have contributed to model performance, its significance is extremely limited and the parameters of the EP methods themselves are more critical determinants.

### v2

@@ Appendix E More experiments
E.1 Effect of numbers of GCN Layers
We explore the impact of GCN depth on accuracy by testing GCNs with 4, 6, and 8 layers, using our edge perturbation methods alongside SPAN baselines. Experiments were conducted with the GRACE and G-BT frameworks on the Cora dataset for node classification and the MUTAG dataset for graph classification. Each configuration was run three times, with the mean accuracy and standard deviation reported.
Overall, deeper GCNs (6 and 8 layers) tend to perform worse across both tasks, reinforcing the observation that deeper architectures, despite their theoretical expressive power, may negatively impact the quality of learned representations. The results are summarized in Tables 7 and 8.
Table 7: Impact of GCN depth on node classification task on the Cora dataset. The best result of each column is in grey. Metric is accuracy (%).
Model 4 6 8
GBT+DropEdge 83.53$\pm$ 1.48 82.06$\pm$ 3.45 80.88$\pm$ 1.38
GBT +AddEdge 81.99$\pm$ 0.79 79.04$\pm$ 1.59 79.41$\pm$ 1.98
GBT+SPAN 80.39$\pm$ 2.17 81.25$\pm$ 1.67 79.41$\pm$ 1.87
GRACE+DropEdge 82.35$\pm$ 1.08 82.47$\pm$ 1.35 81.74$\pm$ 2.42
GRACE +AddEdge 79.17 $\pm$1.35 78.80$\pm$ 0.96 81.00$\pm$ 0.17
GRACE+SPAN 80.15$\pm$ 0.30 80.15$\pm$ 0.79 75.98$\pm$ 1.54
Table 8: Impact of GCN depth on graph classification task on the MUTAG dataset. The best result of each column is in grey. Metric is accuracy (%).
Model 4 6 8
GBT+DropEdge 90.74 $\pm$ 2.61 $8$$\pm$ 4.53 $8$$\pm$ 7.85
GBT +AddEdge 94.44 $\pm$ 0.00 94.44 $\pm$ 4.53 94.44 $\pm$ 4.53
GBT+SPAN 94.44 $\pm$ 4.53 92.59 $\pm$ 2.61 90.74 $\pm$ 2.61
GRACE+DropEdge 94.44 $\pm$ 0.00 90.74 $\pm$ 2.61 90.74 $\pm$ 2.61
GRACE +AddEdge 92.59 $\pm$ 5.23 94.44 $\pm$ 4.53 94.44 $\pm$ 0.00
GRACE+SPAN 90.74 $\pm$ 2.61 90.74 $\pm$ 5.23 $8$$\pm$ 7.85
E.2 Effect of GNN encoder
To further validate the generality of our approach, we conducted additional experiments using different GNN encoders. For the node classification task, we evaluated the Cora dataset with GAT as the encoder, while for the graph classification task, we performed experiments on the MUTAG dataset using both GAT and GPS as encoders.
The results, presented in Tables 9 and 10, are shown alongside the results obtained with GCN encoders. These findings demonstrate that our simple edge perturbation method consistently outperforms the baselines, regardless of the choice of the encoder. This confirms that our conclusions hold across different encoder architectures, underscoring the robustness and effectiveness of the proposed approach.
Table 9: Accuracy of node classification with different GNN encoders on Cora dataset. The best result of each column is in grey. Metric is accuracy (%).
Model GCN GAT
MVGRL+SPAN 84.57 $\pm$ 0.22 82.90 $\pm$ 0.86
MVGRL+DropEdge 84.31 $\pm$ 1.95 83.21 $\pm$ 1.41
MVGRL +AddEdge 83.21 $\pm$ 1.65 83.33 $\pm$ 0.17
GBT+SPAN 82.84 $\pm$ 0.90 83.47 $\pm$ 0.39
GBT + DropEdge 84.19 $\pm$ 2.07 84.06 $\pm$ 1.05
GBT + AddEdge 85.78 $\pm$ 0.62 81.49 $\pm$ 0.45
GRACE + SPAN 82.84 $\pm$ 0.91 82.74 $\pm$ 0.47
GRACE + DropEdge 84.19 $\pm$ 2.07 82.84 $\pm$ 2.58
GRACE + AddEdge 85.78 $\pm$ 0.62 82.84 $\pm$ 1.21
BGRL + SPAN 83.33 $\pm$ 0.45 82.59 $\pm$ 0.79
BGRL + DropEdge 83.21 $\pm$ 3.29 80.88 $\pm$ 1.08
BGRL + AddEdge 81.49 $\pm$ 1.21 82.23 $\pm$ 2.00
Table 10: Accuracy of graph classification with different GNN encoders on MUTAG dataset. The best result of each column is in grey. Metric is accuracy (%).
Model GCN GAT GPS
MVGRL+SPAN 93.33 $\pm$ 2.22 96.29 $\pm$ 2.61 94.44 $\pm$ 0.00
MVGRL+DropEdge 93.33 $\pm$ 2.22 92.22 $\pm$ 3.68 96.26 $\pm$ 5.23
MVGRL +AddEdge 94.44 $\pm$ 3.51 94.44 $\pm$ 6.57 95.00 $\pm$ 5.24
GBT+SPAN 90.00 $\pm$ 6.47 94.44 $\pm$ 4.53 90.74 $\pm$ 5.23
GBT + DropEdge 92.59 $\pm$ 2.61 94.44 $\pm$ 4.53 94.44 $\pm$ 4.53
GBT + AddEdge 92.59 $\pm$ 2.61 92.59 $\pm$ 2.61 94.44 $\pm$ 4.53
GRACE + SPAN 90.00 $\pm$ 4.15 96.29 $\pm$ 2.61 92.59 $\pm$ 2.61
GRACE + DropEdge $8$$\pm$ 3.51 94.44 $\pm$ 0.00 94.44 $\pm$ 4.53
GRACE + AddEdge 92.22 $\pm$ 4.22 96.29 $\pm$ 2.61 94.44 $\pm$ 0.00
BGRL + SPAN 90.00 $\pm$ 4.15 94.44 $\pm$ 4.53 94.44 $\pm$ 0.00
BGRL + DropEdge $8$$\pm$ 4.96 90.74 $\pm$ 4.54 92.59 $\pm$ 5.23
BGRL + AddEdge 91.11 $\pm$ 5.66 96.29 $\pm$ 2.61 96.29 $\pm$ 2.61
E.3 Relationship between spectral cues and performance of EP
@@
Based on the findings obtained from Sec 7.1, it is very likely that spectral information can not be distinguishable enough for good representation learning on the graph. But to more directly answer the question of whether spectral cues and information play an important role in the learning performance of EP, we continue to conduct a statistical analysis to evaluate the influence of various factors on the learning performance. The results turn out to be consistent with our claim that spectral cues are insignificant aspects of outstanding performance on accuracy observed in Sec. 6.
E.3.1 Statistical analyses on key factors on performance of EP
@@
From a statistical angle, we have a few dimensions of factors that can possibly influence learning performance, like the parameters of EP (i.e. drop rate $p$ in DropEdge or add rate $q$ in AddEdge) as well as potential spectral cues lying in the argument graphs. Therefore, to rule out the possibility that spectral cues and information are significant, comparisons are conducted on the impact of the parameters of EP in the augmentations versus:
@@
The polynomial regression analysis in Table 11 highlights that the drop rate $p$ is the primary factor influencing model performance, showing strong and significant linear and non-linear relationships with test accuracy. In contrast, both the OG-AUG and AUG-AUG spectral distances have relatively minor impacts on performance, indicating that they are not significant determinants of the model’s efficacy.
@@
Table 12: IV2SLS regression results for the node-level task. The parameter $p$ refers to the drop rate in DropEdge. The experiment comes in pairs for each pair of variables and the better result is marked in grey.
@@
(X = AUG-AUG, Z = $p$) 0.341 45.77 1.68e-08
(Z = $p$ ,Z = AUG-AUG) 0.611 47.85 9.85e-09
(X = OG-AUG, Z = $p$) 0.250 40.22 7.51e-08
(X = $p$, Z = OG-AUG) 0.606 41.27 5.62e-08
The IV2SLS regression results for the node-level task in Table 12 indicate that the edge perturbation parameters are more significant determinants of model performance than spectral distances. Specifically, when the spectral distance between augmented graphs (AUG-AUG) is the explanatory variable (X) and drop rate $p$ are the instrument (Z), the model explains 34.1% of the variance in performance (R-squared = 0.341). Conversely, when the roles are reversed (X = $p$, Z = AUG-AUG), the model explains 61.1% of the variance (R-squared = 0.611), indicating a stronger influence of edge perturbation parameter $p$. A similar conclusion can be made when comparing OG-AUG and $p$.
@@
The analyses distinctly show that the direct edge perturbation parameters have a consistently stronger and more significant impact on model performance than the two types of spectral distances that serve as a reflection of spectral information. The results support the argument that while spectral information might have contributed to model performance, its significance is extremely limited and the parameters of the EP methods themselves are more critical determinants.